OpenStudy (anonymous):

Please HELP What's the equation of the line perpendicular to 10x + 3y = 5 that passes through (3, 1)?

OpenStudy (anonymous):

First, what is the slope of the line given? now what is the line of the perpendicular line?

OpenStudy (anonymous):

useing this example * Parallel lines have the same slope, e.g. y = 3x - 2 and y = 3x + 500 are parallel, since both lines have the same slope (= 3). * Lines that are perpendicular their slopes are NEGATIVE reciprocals, i.e. they are flipped and have opposite signs, i.e. y = 2/3 x + 7 and y = -3/2 x - 4 are perpendicular, since the slopes are negative reciprocals (i.e. 2/3 vs -3/2). Find the line through the point (1, -3) that is PARALLEL to the line y = 1/4 x - 2. Parallel line implies the SAME slope, i.e. m = 1/4 y - y1 = m(x - x1) <-- point slope formula for a line y - -3 = 1/4 (x - 1) <-- plug in the point and the slope y + 3 = 1/4 x - 1/4 <-- minus a minus = positive; distribute the 1/4 into the parentheses; subtract 3 = 12/4 from both sides y = 1/4 x - 13/4 Find the line through the point (1, -3) that is PERPENDICULAR to the line y = 1/4 x - 2. Perpendicular means the slope flipped and switched, i.e. m = 1/4 --> m = -(4/1) = - 4 y - y1 = m(x - x1) <-- point slope formula for a line y - -3 = - 4 (x - 1) <-- plug in the point and the slope y + 3 = - 4 x + 4 <-- minus a minus = positive; distribute the - 4 into the parentheses; subtract 3 from both sides y = - 4 x + 1

OpenStudy (anonymous):

LOL - use the problem you posted. Rewrite the given problem into y = mx + b M tells you what slope is. Try it again, what is the slope of your given problem

OpenStudy (anonymous):

umm 10

OpenStudy (anonymous):

that is not correct

OpenStudy (anonymous):

y=mx+b move the 10x to the other side by subtracting 10

OpenStudy (anonymous):

you get a negative slope of a 10

OpenStudy (anonymous):

First, take one step at a time. Tell me what the given equation would be when you rewrite to y = mx + b

OpenStudy (anonymous):

Cheetos that is not correct either....

OpenStudy (anonymous):

than you divide everything by 3

OpenStudy (anonymous):